If given 290ml of a diluted glucose solution 0.75% (w/v) that was made by adding 30mL of a stock glucose solution and 260mL of water, what was the original concentration of the stock glucose solution expressed as grams/100mL or % (w/v)?
Answer #1

Given,

260(Water)ml+30ml(glucose stocked=solution(ml)+solvent(g))=290ml(Total Solution)

According to given w/v(Weight in grams/Total solution volume)=0.75 for 290ml
w/290=0.75%=0.75/100
=>w=0.75x290/100=217.5/100
w=2.17g/ml

Weight of glucose in 290ml solution = weight of glucose in 30ml(Because of no added glucose for any of the solutions)

Since we have w=2.17g/100ml(Glucose weight for 290ml & 30ml remains same) apply formula for 30ml

%=2.17(w)*100/30(v)=7.25%


Please observe the colouring, so you may not have any doubts further in using formula.